### v1

# Latent-space variational data assimilation in two-dimensional turbulence

Abstract
Starting from limited measurements of a turbulent flow, data assimilation (DA) attempts to estimate all the spatio-temporal scales of motion. Success is dependent on whether the system is observable from the measurements, or how much of the initial turbulent field is encoded in the available measurements. Adjoint-variational DA minimises the discrepancy between the true and estimated measurements by optimising the initial velocity or vorticity field (the ‘state space’). Here we propose to instead optimise in a lower-dimensional latent space which is learned by implicit rank minimising autoencoders. Assimilating in latent space, rather than state space, redefines the observability of the measurements and identifies the physically meaningful perturbation directions which matter most for accurate prediction of the flow evolution. When observing coarse-grained measurements of two-dimensional Kolmogorov flow at moderate Reynolds numbers, the proposed latent-space DA approach estimates the full turbulent state with a relative error improvement of two orders of magnitude over the standard state-space DA approach. The small scales of the estimated turbulent field are predicted more faithfully with latent-space DA, greatly reducing erroneous small-scale velocities typically introduced by state-space DA. These findings demonstrate that the observability of the system from available data can be greatly improved when turbulent measurements are assimilated in the right space, or coordinates.
1 Introduction
Assimilating experimental data into numerical simulations improves the fidelity of the simulations and enables nonintrusive access to all the scales of the estimated flow. However, the estimation of turbulence from limited measurements is a difficult ill-posed problem (zaki2025). Turbulence presents challenges such as the chaotic nature of the forward and dual problems, the non-uniqueness of solutions consistent with the measurements, and the introduction of erroneous small-scale velocities that decay over the solution trajectory.
Conventional data assimilation utilizes the measurements to navigate the state-space representation of turbulence, and to directly estimate the velocity or vorticity field that justifies the measurements. However, the state-space representation of turbulence is not necessarily suitable for this task. In this work, we propose to first map from the measurements to a pre-designed latent space, from which the full turbulent field can be subsequently decoded. We show that the accuracy of the estimated turbulent field can be significantly improved over a broad range of scales by interpreting the turbulent measurements in this latent space, when compared to the estimation using the state-space coordinates.
The term latent space refers to a low-dimensional and interpretable representation of the turbulent field, whether by familiar modal decompositions (taira2017modal), or by non-linear autoencoder transformations (brunton2020). Physical insights of complex dynamical systems can be gleaned from these low-dimensional latent spaces (Fukami2023). Rank-minimising autoencoders have successfully learned parsimonious representations of chaotic systems such as the Lorenz system, the Kuramoto-Sivashinsky equation and the lambda-omega reaction-diffusion system (Zeng_2024). The latent spaces of such autoencoders have yielded insights on the nature of bursting events and the dynamical relevance of unstable periodic orbits in forced two-dimensional (2D) turbulence (cleary2025).
In the context of turbulence estimation, data-driven methods have been used to directly super-resolve limited instantaneous observations to the full flow state (e.g. Fukami_Fukagata_Taira_2019). Another super-resolution study considered a time-history of scarce measurements and inferred the pressure field of forced isotropic turbulent flow (williams2024). Machine-learning and classical DA methods to estimate turbulence have been pursued mostly separately, with a few exceptions including the following examples: DU2023 compared the estimation of wall turbulence using physics-informed neural networks and adjoint-variational techniques. Page_2025 modified the training of super-resolution networks to incorporate a time-forward Navier-Stokes evolution in the output, and a comparison to future data. Most recently, (weyrauch2025) used the output of this super-resolution network as an initial guess to adjoint-variational DA. While this last effort has interfaced the data-driven methods and adjoint-variational DA, the two techniques were not fully integrated.
In the present work, we integrate the two turbulence estimation strategies. We exploit a learned space that is discovered using data-driven methods and the physics constraints of adjoint-variational DA, to significantly enhance the observability of turbulence systems and their estimation across a wider range of scales. In §2, we outline the flow configuration studied and summarise the standard variational DA procedure. In §3, we present the proposed latent-space DA procedure. In §4, we compare and interpret the improved performance of latent-space DA. We conclude in §5.
2 State-space data assimilation
We consider Kolmogorov flow, which is monochromatically forced 2D turbulence on a square and doubly periodic domain (Chandler2013). The out-of-plane vorticity $\omega=\partial_{x}v-\partial_{y}u$ defines the flow state in state space and evolves according to
[…]
Figure 1: Schematic of latent-space data assimilation. The latent representation $\eta$ is mapped to state space $\omega$ by the decoder $\mathcal{F}_{\mathcal{D}}$, where the adjoint field $\omega^{\dagger}$ is computed. The grey arrow closes the standard, state-space ‘loop’. In latent-space assimilation, the latent state is updated by the transformed adjoint field $\eta^{\dagger}$. Sensor resolution is indicated by the lattice of black dots in the lower left corner of $\omega$.
Our objective is to estimate the flow state $\omega^{*}_{0}$ which, when evolved using the Navier-Stokes equations (1), reproduces available measurements $m_{n}^{R}=\mathcal{M}(\omega^{R}_{n})\in\mathbb{R}^{d_{m}}$ from a reference solution $\omega^{R}$ at discrete times $t_{n}=n\Delta t$ over the time horizon $t_{n}\in[0,T]$ for $n=0,\dots,N$. The problem is formulated as a variational minimisation of a cost function of the discrepancy between the estimated and true measurements,
$\mathcal{J}(\omega_{0})=\frac{1}{2}\sum_{n=0}^{N}\left\|\mathcal{M}(\omega_{n})-m^{R}_{n}\right\|^{2},$ (2)
subject to the constraint that $\omega_{n}=f^{t_{n}}(\omega_{0})$ is the time-forward map of (1) from initial condition $\omega_{0}$. The required gradient of (2) with respect to $\omega_{0}$ can be computed using the discrete adjoint (wang2019discrete) or automatic differentiation (fan2025).
The measurement operator $\mathcal{M}$ is defined to be the coarse-graining operation $\mathcal{M}:\mathbb{R}^{N_{x}\times N_{y}}\to\mathbb{R}^{N_{x}/M\times N_{y}/M}$ which samples the high-resolution data at every $M^{\text{th}}$ gridpoint in both $x$- and $y$-directions. The temporal coarsening $\Delta t=M\delta t$ is set by the same coarsening factor, where $\delta t$ is the time-step of the numerical simulation. At $Re=\{40,100\}$, coarsening is set to $M=16$ and to $M=32$ at $Re=400$. The DA time horizon is $T\approx 0.6T_{L}$ where $T_{L}$ is the Lyapunov timescale at each $Re$.
As represented by the blue box in figure 1, the gradient of (2) can be computed by solving the adjoint equations,
$\frac{\partial\omega^{\dagger}}{\partial\tau}+J(\psi,\omega^{\dagger})-\frac{1}{Re}\nabla^{2}\omega^{\dagger}+\psi^{\dagger}=\frac{\mathcal{D}\mathcal{J}}{\mathcal{D}\omega}\,,\qquad\nabla^{2}\psi^{\dagger}-J(\omega,\omega^{\dagger})=\frac{\mathcal{D}\mathcal{J}}{\mathcal{D}\psi}\,$, (3)
backwards in time, where $\tau:=T-t$, the streamfunction $\psi$ is related to the velocity components via $u=\partial_{y}\psi,v=-\partial_{x}\psi$, and $J(\psi,\omega)=\frac{\partial\psi}{\partial x}\frac{\partial\omega}{\partial y}-\frac{\partial\psi}{\partial y}\frac{\partial\omega}{\partial x}$. The adjoint field $\omega^{\dagger}$ at $t=0$ yields the variation of the cost function with respect to the initial condition $\omega_{0}$,
$\frac{\mathscr{D}\mathcal{J}}{\mathscr{D}\omega_{0}}=\omega^{\dagger}(t=0),$ (4)
which can be used to update the estimated $\omega_{0}$ in a direction that minimizes the cost by better reproducing the measurements. In lieu of explicit solution of the adjoint equations (3), we compute the gradient (4) by automatic differentiation. The cost function is then minimised using the Adam optimiser for a total of 500 optimisation steps at each $Re$. A sweep over initial step sizes (or learning rates) was performed, and $\alpha=0.2$ was selected.
Variational DA in state space requires a first guess of the initial flow field $\omega_{0}$. We will consider two initialisation approaches. The first is a bicubic interpolation of the measurements (InterpDA). The second is using a pre-trained super-resolution (SR) network $\mathcal{F}_{SR}:\mathbb{R}^{d_{m}}\to\mathbb{R}^{d_{\omega}}$, which maps from instantaneous measurements to instantaneous full resolution fields. This initialisation for data assimilation (SR-DA) was shown to yield significant improvement in the accuracy of the assimilation estimate by weyrauch2025. We therefore adopt their SR network for comparison. These assimilations in state space serve as a benchmark for the proposed latent-space assimilation.
3 Latent-space data assimilation
The classical variational-DA algorithm attempts to identify the optimal initial vorticity, or state-space representation of the flow, to reproduce the measurement. The main idea of our latent-space DA is to, instead, attempt to predict a latent-space representation $\eta\in\mathbb{R}^{d_{\eta}}$ (see the orange box in figure 1). Shifting the object of the optimisation to the latent space requires an evaluation of the gradient with respect to the new latent coordinates. The latent space is mapped to state space by a pre-trained decoder $\mathcal{F}_{\mathcal{D}}(\eta)=\omega\in\mathbb{R}^{d_{\omega}}$. The new DA cost function is then
$\mathcal{J}(\eta_{0})=\frac{1}{2}\sum_{n=0}^{N}\left\|\left[\mathcal{M}\circ f^{t_{n}}\circ\mathcal{F}_{\mathcal{D}}\right](\eta_{0})-m^{R}_{n}\right\|^{2},$ (5)
where the latent state estimate $\eta_{0}$ is first decoded to state space, then (1) is solved in state space and finally the model observations are compared to available measurements (combined orange and blue boxes in figure 1). The adjoints in latent and state space are related by the Jacobian $\partial\mathcal{F}_{\mathcal{D}}/\partial\eta\in\mathbb{R}^{d_{\omega}\times d_{\eta}}$,
$\eta^{\dagger}=\left(\frac{\partial\mathcal{F}_{\mathcal{D}}}{\partial\eta}\right)^{\top}\omega^{\dagger}.$ (6)
We note that no dynamical model in the latent space is required and all time marching is performed in state space, such that the estimated solution $f^{t}(\mathcal{F}_{\mathcal{D}}(\eta_{0}))$ for $t\in[0,T]$ exactly satisfies the Navier-Stokes equations (1). Furthermore, given a pre-existing adjoint solver for the flow in question, only the ability to compute vector-Jacobian products of the decoder is required to perform latent-space DA according to (6). As in state-space DA, the Adam optimiser is used to minimise (5), and the same number of optimisation steps were taken. A sweep over initial step sizes led to the choices $\alpha=\{5,0.2,0.01\}$ at $Re=\{40,100,400\}$.
3.1 The latent space
We adopt the latent space of the implicit rank-minimising autoencoder (IRMAE). Unlike standard autoencoders, IRMAE is distinguished by the addition of a series of fully-connected, linear layers in the bottleneck of the network, which drive down the intrinsic dimensionality of the latent representation (jing2020; Zeng_2024). The IRMAE architecture has demonstrated comparable performance to variational autoencoders for smooth interpolation in the latent space and generating new samples from random noise (jing2020). As such, the latent representation exhibits these favourable properties in addition to being approximately minimal rank.
The IRMAE network $\mathscr{A}$ seeks to learn the identity function
$\mathscr{A}(\omega)\equiv[\mathcal{F}_{\mathcal{D}}\circ\mathcal{W}\circ\mathcal{F}_{\mathcal{E}}](\omega)\approx\omega,$ (7)
where the encoder $\mathcal{F}_{\mathcal{E}}:\mathbb{R}^{d_{\omega}}\to\mathbb{R}^{d_{\eta}}$ maps the input vorticity snapshot to a low-dimensional representation ($d_{\eta}=1024$), $\mathcal{W}:\mathbb{R}^{d_{\eta}}\to\mathbb{R}^{d_{\eta}}$ represents a series of four fully-connected, equally-sized linear layers (pure matrix multiplication) within the embedding space, and the decoder $\mathcal{F}_{\mathcal{D}}$ is defined as above. The encoder and decoder consist of a series of convolutional dense blocks at varying resolutions.
To train the networks at $Re=\{40,100,400\}$, datasets were generated by sampling long-time trajectories of the flow at every time unit, resulting in datasets with a total number of snapshots $N_{S}\approx\{6,10,10\}\times 10^{4}$, respectively. The networks are then trained to minimise the loss function
$\mathscr{L}=\frac{1}{N_{S}}\sum_{j=1}^{N_{S}}\left\|\mathscr{A}(\omega^{j})-\omega^{j}\right\|^{2},$ (8)
where each $\omega^{j}$ is a full-field snapshot from the training dataset. Full details of the IRMAE architecture, training protocol and reconstruction accuracy of (7) are provided in the supplementary material. For each $Re$ considered, DA was performed on independent trajectories which were never seen during the training of the neural networks.
Variational DA in latent space requires an first guess of the initial latent representation $\eta_{0}$. This initialisation is obtained by using the aforementioned pre-trained SR network followed by the IRMAE encoder, i.e., $\eta_{0}=\mathcal{F}_{\mathcal{E}}\circ\mathcal{F}_{SR}(m^{R}_{0})$. Comparable performance can be achieved by training a fully-connected network to map directly from measurement space to the IRMAE latent space.
4 Results
4.1 Accuracy of the estimated fields
Figure 2: Relative error $\varepsilon$ of the estimated turbulent field by (blue) InterpDA, (orange) SR, (green) SR-DA and (red) LatentDA at (left) $Re=40$ with $M=16$, (middle) $Re=100$ with $M=16$ and (right) $Re=400$ with $M=32$ as a function of time normalised by the Lyapunov timescale. Bold lines: ensemble average of ten independent trajectories; Coloured regions: max/min values attained by the ensemble. Grey region marks the data-assimilation time horizon.
The accuracy of the flow trajectories is compared when the initial flow states are estimated from super-resolution (SR), super-resolution-initialised DA in state space (SR-DA), and DA in latent space (LatentDA). Standard state-space DA initialised with a bicubic interpolation (InterpDA) is presented as a benchmark. The evolution of the error,
$\varepsilon(t)={\|f^{t}(\omega^{*}_{0})-f^{t}(\omega^{R}_{0})\|}\,/\,{\|f^{t}(\omega^{R}_{0})\|},$ (9)
for each method at $Re=\{40,100,400\}$ is reported in figure 2, where $\omega^{*}_{0}$ and $\omega^{R}_{0}$ are the estimated and reference turbulent fields at $t=0$. The spatio-temporal coarsening factor was $M=16$ at both $Re=\{40,100\}$ and $M=32$ at $Re=400$, and DA was performed over the time horizon $T\approx 0.6T_{L}$ which is indicated by the grey shaded region in figure 2.
Figure 3: Comparison of (a) reference $\omega^{R}_{0}$ at $t=0$ and the estimated field $\omega^{*}_{0}$ from (i-iv) InterpDA, SR, SR-DA, LatentDA at $Re=400$. (b) Contours of the out-of-plane vorticity, (c) enstrophy spectra $\Omega(k)$ as a function of wavenumber $k$. The black spectra and the black contours in (biv) denote the reference data.
LatentDA (red) results in an improvement of approximately two orders of magnitude over the standard InterpDA approach (blue) and one order of magnitude over SR-DA (green) at $Re=\{40,100\}$. Even at $Re=400$, we still observe an improvement by more than one order of magnitude over the InterpDA approach, and a significant improvement over the best-performing approach in state space (SR-DA). Since the accuracy advantage of the estimated trajectories from LatentDA is retained throughout the observation horizon, and because the growth rate of $\varepsilon(t)$ after $t=T$ is consistent across all three DA approaches, accurate predictions from LatentDA can be made over much longer time intervals. For example, at $Re=100$ and using LatentDA, the prediction error is ${\sim}1\%$ at $2.5T_{L}$ (not shown), which is a four folds longer horizon than when using SR-DA.
A comparison of the estimated $\omega^{*}_{0}$ fields at $Re=400$ using each approach is presented in figure 3. When the data is assimilated in state space (figure 3(bi)), $\omega^{*}_{0}$ exhibits unphysical high-wavenumber artefacts. In contrast, the estimated fields by SR (figure 3(bii)) is overly smooth, which is a symptom of the spectral bias of neural networks. Assimilating this field (SA-DA) introduces some high-frequency errors, but the enstrophy spectrum $\Omega(k)$ remains poorly representative of the true field. When the assimilation is performed in latent space, the predicted initial field is most accurate and $\Omega(k)$ is most consistent with the reference turbulent state. Similar trends were observed at the lower $Re$ considered.
Figure 4: (a) The spectra $\Omega_{adj}$ of the (green) adjoint field in state space $\omega^{\dagger}$ and (red) latent adjoint field decoded to state space $\mathcal{F}_{\mathcal{D}}(\eta+\alpha\eta^{\dagger})-\mathcal{F}_{\mathcal{D}}(\eta)$. Contours of the (bi) state-space adjoint and (bii) decoded latent adjoint fields.
To understand the cause of the high-wavenumber artefacts in state-space DA, it should be noted that the adjoint equations (3) are forced by $\mathcal{D}\mathcal{J}/\mathcal{D}\omega$, which is a series of singular impulses in space and time at each measurement location. These superpositions of delta functions are advected and diffused by (3), but signatures of this forcing are apparent in the visualisation of $\omega^{\dagger}(t=0)$ in figure 4(bi). As the Fourier transform of a delta function is a constant function of $k$, these singular impulses lead to high wavenumber spectral content of the state-space adjoint variables (figure 4(a)). Each variational DA iteration slightly perturbs the turbulent field estimate in a direction with high energy in the high wavenumbers. For latent-DA, the effective adjoint update can be visualised in state space by evaluating $\mathcal{F}_{\mathcal{D}}(\eta+\alpha\eta^{\dagger})-\mathcal{F}_{\mathcal{D}}(\eta)$. As shown in figure 4, the energy in the high wavenumbers of this effective latent adjoint direction is reduced to machine precision, which is consistent with the improved spectrum of $\omega_{0}^{*}$ for LatentDA.
4.2 Observability in latent space
We consider two dynamical perspectives that explain the improved performance of LatentDA. The first perspective recalls that the IRMAE models have been trained on data sampled from the turbulent attractor $\omega\in\mathcal{A}$ and have learned latent representations $\eta$ such that $\mathcal{F}_{\mathcal{D}}(\eta)\in\mathcal{A}$ holds approximately. Let us consider the gradient direction in latent space, $\alpha\eta^{\dagger}$, as a small perturbation such that $\mathcal{F}_{\mathcal{D}}(\eta-\alpha\eta^{\dagger})\in\mathcal{A}$ still holds approximately. By expanding this small perturbation to linear order,
$\mathcal{F}_{\mathcal{D}}(\eta-\alpha\eta^{\dagger})$ $=\mathcal{F}_{\mathcal{D}}(\eta)-\alpha\frac{\partial\mathcal{F}_{\mathcal{D}}}{\partial\eta}\eta^{\dagger}+\dots\in\mathcal{A},$
it is clear that the columns of the decoder Jacobian are the perturbation directions which, to linear order, approximately remain on the turbulent attractor. The linear transformation of adjoints from state to latent space (6) can now be understood as a projection onto these physically relevant perturbation directions. The associated linearised update in state space is given by,
$\omega_{0}^{*}\approx\mathcal{F}_{\mathcal{D}}(\eta)-\alpha\frac{\partial\mathcal{F}_{\mathcal{D}}}{\partial\eta}\left(\frac{\partial\mathcal{F}_{\mathcal{D}}}{\partial\eta}\right)^{\top}q^{\dagger}.$
As such, the estimate remains physically relevant in state space throughout the latent variational DA method, to linear order. Physically, these improved perturbation directions are marked by an absence of the high-wavenumber artefacts discussed previously.
The second perspective considers how observable the reference initial turbulent field is from the measurements. As we will show here, the iterative gradient-based updates throughout the variational DA method are also expansions in some basis of adjoint fields. When assimilating data variationally, it is beneficial for the reference turbulent field $\omega^{R}_{0}$ to be well represented in this basis, such that the iterative updates can more effectively converge onto $\omega^{R}_{0}$. To begin, we define the deviation field $w=\omega-\omega^{R}$ which is governed by the linearised Navier-Stokes equations
$\partial_{t}w-J(\varphi,\omega^{R})-J(\psi^{R},w)=\frac{1}{Re}\nabla^{2}w,$ (10)
with the associated deviation streamfunction $\varphi=-\nabla^{2}w$ and the ground truth streamfunction $\psi^{R}=-\nabla^{2}\omega^{R}$. As shown in wang2022observable for turbulent channel flow, the DA cost function (2) can be written in terms of the measurement kernel $\phi(\bm{x}_{m})$ which extracts the measurement of interest at location $\bm{x}_{m}$
$\mathcal{J}(w_{0})=\frac{1}{2}\sum_{n=0}^{N}\sum_{m=1}^{d_{m}}\left\langle w(t=t_{n}),\phi(\bm{x}_{m})\right\rangle^{2},$ (11)
where $\langle a,b\rangle=\int_{V}ab\,dV$ is the spatial inner product over the computational domain $V$. In this study $\phi(\bm{x}_{m})=\delta(\bm{x}-\bm{x}_{m})$ is the Kronecker delta function, but more complex kernels can be defined as required. Forward-adjoint duality can be exploited to write (11) in terms of the adjoint field $w^{\ddagger}$, as
$\left\langle w(t=t_{n}),\phi(\bm{x}_{m})\right\rangle=\left\langle\mathcal{L}w_{0},\phi(\bm{x}_{m})\right\rangle=\left\langle w_{0},\mathcal{L}^{\ddagger}\phi(\bm{x}_{m})\right\rangle=\left\langle w_{0},w^{\ddagger}(t=0;t=t_{n},\bm{x}_{m})\right\rangle,$
where $\mathcal{L}$ is the forward operator of the linearised Navier-Stokes equations (10) which advances $w_{0}$ to $w(t=t_{n})$. Note that $w^{\ddagger}\neq\omega^{\dagger}$, and an explicit relation between them will be given below. The associated adjoint operator $\mathcal{L}^{\ddagger}$ solves the linearised adjoint equations
$\frac{\partial w^{\ddagger}}{\partial\tau}+J(\psi^{R},w^{\ddagger})-\frac{1}{Re}\nabla^{2}w^{\ddagger}+\varphi^{\ddagger}=0\,,\quad\nabla^{2}\varphi^{\ddagger}-J(\omega^{R},w^{\ddagger})=0\,$, (12)
with the initial condition $w^{\ddagger}(\tau=0)=\phi(\bm{x}_{m})$. This duality can be used to rewrite (11) as
$\mathcal{J}(w_{0})=\frac{1}{2}\sum_{n=0}^{N}\sum_{m=1}^{d_{m}}\left\langle w_{0},w^{\ddagger}(t=0;t=t_{n},\bm{x}_{m})\right\rangle^{2},$ (13)
enabling the explicit expression of the gradient,
$\omega^{\dagger}(t=0)=\frac{\mathscr{D}\mathcal{J}}{\mathscr{D}w_{0}}=\sum_{n=0}^{N}\sum_{m=1}^{d_{m}}\left\langle w_{0},w^{\ddagger}(t=0;t=t_{n},\bm{x}_{m})\right\rangle w^{\ddagger}(t=0;t=t_{n},\bm{x}_{m}),$ (14)
and Hessian of the DA cost function,
$\mathcal{H}\coloneq\frac{\mathscr{D}^{2}\mathcal{J}}{\mathscr{D}w_{0}\mathscr{D}w_{0}}=\sum_{n=0}^{N}\sum_{m=1}^{d_{m}}w^{\ddagger}(t=0;t=t_{n},\bm{x}_{m})w^{\ddagger}(t=0;t=t_{n},\bm{x}_{m}).$ (15)
The gradient (14) is then an expansion in the basis spanned by the adjoint fields $w^{\ddagger}(t=0;t=t_{n},\bm{x}_{m})$, while the Hessian (15) can be written as the cross-correlation $\mathcal{H}=AA^{\top}$ of the matrix $A$ with these adjoint fields as columns. In state space, $\mathcal{H}$ is computed at the estimated $\omega^{*}_{0}$, and in latent space $\mathcal{H}=\mathscr{D}^{2}\mathcal{J}/\mathscr{D}\eta_{0}\mathscr{D}\eta_{0}$ is computed at the estimated $\eta^{*}_{0}$.
Figure 5: (a) The relative error $\varepsilon$ of the instantaneous reconstruction of the reference turbulent field $\omega^{R}_{0,i}$ (lines) and the time-averaged relative error over the DA time horizon (circles) using the first $i$ adjoint POD modes in (green) state and (red) latent space at $Re=100$. (b) Contours of the out-of-plane vorticity of the reconstructed field with 500 adjoint POD modes in (i) state and (ii) latent space, and (iii) the reference turbulent field.
The eigendecomposition $\mathcal{H}v_{i}=\lambda_{i}v_{i}$ with eigenvalues $\lambda_{i}$ and eigenvectors $v_{i}$ then yields a proper orthogonal decomposition (POD) basis for these adjoint fields. We examine how well the ground truth turbulent state $\omega^{R}_{0}$ can be represented in this adjoint basis by computing its reconstruction at $t=0$ using the first $i$ POD modes of the Hessian at $\omega^{*}_{0}$,
$\omega^{R}_{0,i}=\sum_{j=0}^{i}\left\langle v_{j},\omega^{R}_{0}\right\rangle v_{j}.$ (16)
The relative error $\varepsilon$ of the reconstruction is reported in figure 5, for $Re=100$, where we compare SR-DA (green) and $\eta_{0}^{*}$ obtained with LatentDA (red). The first 500 adjoint POD modes were computed by the Arnoldi iteration. The reconstructions were also time-evolved over the DA time horizon, and the time-averaged relative errors (circles) were computed as a function of the number of adjoint POD modes used. The reference turbulent state can be reconstructed to ${\sim}5\%$ relative error in the latent adjoint basis with 500 adjoint POD modes, as opposed to ${\sim}50\%$ relative error in the state-space adjoint basis, demonstrating the improved observability of $\omega^{R}_{0}$ from the measurements in the latent space near optimality. Perhaps more notably, ${\sim}100$ modes in latent space can reconstruct the reference turbulent state with ${\sim}20\%$ error, while the same number of modes can only reconstruct the state with ${\sim}75\%$ error in state space. High wavenumber artefacts and signatures of the localized adjoint forcing are evident in the state-space adjoint basis reconstruction (figure 5(bi)). These small-scale fluctuations decay over the DA time horizon, explaining the improved time-averaged relative errors in state space. In contrast, the decoded reconstruction in latent space very closely resembles the reference turbulent state (panels bii-biii), and the error in the representation of the initial state and the time-averaged errors during the evolution are similar. These results demonstrate that the observability of the turbulence using limited measurements can be appreciably improved when the assimilation is performed in the right space, or coordinates.
5 Conclusion
In the study of turbulence, we often rely on the interpretation of measurements to probe the dynamics of the underlying flow. Data assimilation seeks to map from the measurements to the associated turbulent state that satisfies the Navier-Stokes equations. We ask the question if our ability to observe the turbulence can be significantly improved by first mapping from the measurements to a pre-designed latent space, and subsequently to the full turbulent field? We demonstrate that the mapping to the latent intermediate coordinates, namely the latent space of a pre-trained autoencoder, can lead to significant accuracy improvement in the interpretation of turbulence measurements.
In state-space DA, the adjoint field is highly localised in the vicinity of the delta-function forcing, such that the spectrum of the adjoint field is broadband. The updates in latent space are more targetted and physically relevant, resulting in more accurate reconstruction of both the large and small scales, and a smaller departure from the truth over time. An improvement of an order of magnitude was achieved when assimilating data in latent space at $Re=40$ and 100, compared to the best approach considered in state space. A lower but nonetheless appreciable improvement was achieved at $Re=400$, where the increased spatiotemporal complexity of the flow requires a larger latent space dimensionality. The small scales of the estimated flow state are more dynamically relevant at all $Re$ considered.
This work demonstrates the potential benefits of combining variational and data-driven techniques to interpret turbulence measurements. The observability of turbulence from the data is much improved by taking advantage of the latent space of the autoencoder. More generally, the results demonstrate the power of exploiting such new latent spaces in the study of turbulence.
{bmhead}
[Declaration of interests] The authors report no conflict of interest.
{bmhead}
[Supplementary material] Supplementary material is available at …